# Actions Speak Louder than Words: Trillion-Parameter Sequential Transducers for Generative Recommendations

### v1

@@ Appendix B Synthetic Data
The results can be found in Table 1. We always ablate $\text{rab}^{p,t}$ for HSTU as this dataset does not have timestamps. We observe HSTU increasing Hit Rate@10 by more than 100% relative to standard Transformers. Importantly, replacing HSTU’s pointwise attention mechanism with softmax (“HSTU w/ Softmax”) also leads to significant reduction in hit rate, verifying the importance of pointwise attention-like aggregation mechanisms.
Appendix C DLRM Baselines
We give a high level description of the DLRM baselines used in Section 4.
Ranking Setting. The baseline ranking model, as described in (Mudigere et al., 2022), employs approximately one thousand dense features and fifty sparse features. We incorporated various modeling techniques such as Mixture of Experts (Ma et al., 2018), Deep & Cross Network (Wang et al., 2021), target-aware pairwise attention (Zhou et al., 2018), and residual connection over special interaction layers (Zhang et al., 2022).
@@ Appendix C DLRM Baselines
Appendix D Stochastic Length
D.1 Selection of Sampling Technique $\Gamma$
In Equation 4, we introduced a function $\Gamma(S,L)$ which samples a sequence $S$ to length $L$ in order to increase sparsity. Our empirical results indicate that properly designing $\Gamma$ can improve model quality. During the selection of $\Gamma$, we compute a metric $f_{i}=t_{n}-t_{i}$ which corresponds to the amount of time elapsed since the user interacted with item $x_{i}$. We conduct offline experiments with the following samplers:
  • Greedy Sampler – Samples $L$ items with smallest values of $f_{i}$ from $S$
  • Random Sampler – Samples $L$ items from $S$ randomly
  • Feature-Weighted Sampler – Samples $L$ items from $S$ according to a weighted distribution $1-f_{n,i}/(\sum_{j=1}^{L}f_{j,i})$
During our offline experiments, the feature-weighted sampler resulted in the best model quality, as shown in Table 8.
Metric Name Sampler Type
@@ D.1 Selection of Sampling Technique $\Gamma$
Main Engagement Metric 0.495 0.494 0.495
Main Consumption Metric 0.792 0.789 0.791
Table 8: Impact of Sampler $\Gamma(S,L)$ on model quality, measured by Normalized Entropy (NE).
D.2 Impact of Stochastic Length on Sequence Sparsity
@@
D.3 Comparisons Against Sequence Length Extrapolation Techniques
We conduct additional studies to verify that Stochastic Length is competitive against existing techniques for sequence length extrapolation used in language modeling. Many existing methods perform sequence length extrapolation through modifications of RoPE (Su et al., 2023). To compare against existing methods, we train a HSTU variant (HSTU-RoPE) with no relative attention bias and rotary embeddings.
@@ D.3 Comparisons Against Sequence Length Extrapolation Techniques
Appendix E Evaluation of Inference Throughput: GRs w/ M-FALCON vs DLRMs
@@
As discussed in Section 3.4, M-FALCON handles $b_{m}$ candidates in parallel to amortize computation costs across all candidates at inference time. To understand our design, we compare the throughput (i.e., number of candidates scored per second) of GRs and DLRMs based on the same hardware setups. As shown in Figure 9 and Figure 10, GRs’ throughput scales in a sublinear way based on the number of ranking-stage candidates ($m$), up to a certain region – $m=2048$ in our case study – due to cost amortization. Due to attention complexity scaling as $O((n+b_{m})^{2})$, leveraging multiple microbatches by itself improves throughput. Caching further eliminates redundant linear and attention computations on top of microbatching. The two combined resulted in up to 1.65x speedups relative to the $m=1024$ baseline using a single microbatch, as shown in Figure 10. Overall, utilizing M-FALCON, HSTU-based GRs outperform DLRMs in terms of throughput on a large-scale production setup by up to 2.48x, despite GRs being 285x more complex in terms of FLOPs.
@@
Figure 9: End-to-end inference throughtput: DLRMs vs GRs (w/ M-FALCON) in large-scale industrial settings.
@@
Figure 10: End-to-end inference throughtput: M-FALCON throughput scaling, on top of the 285x FLOPs GR model, in large batch settings where $m$ (total number of candidates) ranges from 1024 to 16384 and $b_{m}=1024$.

### v2

@@ Appendix C Evaluation: Synthetic Data
The results can be found in Table 2. We always ablate $\text{rab}^{p,t}$ for HSTU as this dataset does not have timestamps. We observe HSTU increasing Hit Rate@10 by more than 100% relative to standard Transformers. Importantly, replacing HSTU’s pointwise attention mechanism with softmax (“HSTU w/ Softmax”) also leads to a significant reduction in hit rate, verifying the importance of pointwise attention-like aggregation mechanisms.
Appendix D Evaluation: Traditional Sequential Recommender Settings
Our evaluations in Section 4.1.1 focused on comparing HSTU with a state-of-the-art Transformer baseline, SASRec, utilizing latest training recipe. In this section, we further consider two other alternative approaches.
Recurrent neural networks (RNNs). We consider the classical work on sequential recommender, GRU4Rec (Hidasi et al., 2016), to help readers understand how self-attention models, including Transformers and HSTU, compare to traditional RNNs, when all the latest modeling and training improvements are fully incorporated.
Self-supervised sequential approaches. We consider the most popular work, BERT4Rec (Sun et al., 2019), to understand how bidirectional self-supervision (leveraged in BERT4Rec via a Cloze objective) compares with unidirectional causal autoregressive settings, such as SASRec and HSTU.
Method HR@10 HR@50 HR@200 NDCG@10 NDCG@200
ML-1M SASRec (2023) .2828 .5508 .7522 .1545 .2441
BERT4Rec .2500 (-11.6%) .1341 (-13.2%)
GRU4Rec .2811 (-0.6%) .1537 (-0.5%)
ML-20M SASRec (2023) .2906 .5499 .7655 .1621 .2521
BERT4Rec .2816 (-3.4%) .1703 (+5.1%)
GRU4Rec .2813 (-3.2%) .1730 (+6.7%)
Books SASRec (2023) .0292 .0729 .1400 .0156 .0350
Table 12: Evaluations of methods on public datasets in traditional sequential recommender settings (multi-pass, full-shuffle). Compared with Table 4, two other baselines (GRU4Rec and BERT4Rec) are included for completeness.
Results are presented in Table 12. We reuse BERT4Rec results and GRU4Rec results on ML-1M and ML-20M as reported by Klenitskiy & Vasilev (2023). Given a sampled softmax loss is used, we hold the number of negatives used constant (128 for ML-1M, ML-20M and 512 for Amazon Books) to ensure a fair comparison between methods.
The results confirm that SASRec remains one of the most competitive approaches in traditional sequential recommendation settings when sampled softmax loss is used (Zhai et al., 2023a; Klenitskiy & Vasilev, 2023), while HSTU significantly outperforms evaluated transformers, RNNs, and self-supervised bidirectional transformers.
Appendix E Evaluation: Traditional DLRM Baselines
The DLRM baseline configurations used in Section 4 reflect continued iterations of hundreds of researchers and engineers over multiple years and a close approximation of production configurations on a large internet platform with billions of daily active users before HSTUs/GRs were deployed. We give a high level description of the models used below.
Ranking Setting. The baseline ranking model, as described in (Mudigere et al., 2022), employs approximately one thousand dense features and fifty sparse features. We incorporated various modeling techniques such as Mixture of Experts (Ma et al., 2018), variants of Deep & Cross Network (Wang et al., 2021), various sequential recommendation modules including target-aware pairwise attention (one commonly used variant in industrial settings can be found in (Zhou et al., 2018)), and residual connection over special interaction layers (He et al., 2015; Zhang et al., 2022). For the low FLOPs regime in the scaling law section (Section 4.3.1), some modules with high computational costs were simplified and/or replaced with other state-of-the-art variants like DCNs to achieve desired FLOPs.
While we cannot disclose the exact settings due to confidentiality considerations, to the best of our knowledge, our baseline represents one of the best known DLRM approaches when recent research are fully incorporated. To validate this claim and to facilitate readers’ understanding, we report a typical setup based on identical features but only utilizing major published results including DIN (Zhou et al., 2018), DCN (Wang et al., 2021), and MMoE (Ma et al., 2018) (“DLRM (DIN+DCN)”) in Table 7, with the combined architecture illustrated in Figure 9. This setup significantly underperformed our production DLRM setup by 0.71% in NE for the main E-Task and 0.57% in NE for the main C-Task (where 0.1% NE is significant).
Figure 9: A high level architecture of a baseline DLRM ranking model (“DLRM (DIN+DCN)” in Table 7) that utilizes major published work including DIN (Zhou et al., 2018), DCN (Wang et al., 2021), and MMoE (Ma et al., 2018).
@@ Appendix E Evaluation: Traditional DLRM Baselines
Features and Sequence Length. The features used in both of the DLRM baselines, including main user interaction history that is utilized by various sequential encoder/pairwise attention modules, are strict supersets of the features used in all GR candidates. This applies to all studies conducted in this paper, including those used in the scaling studies (Section 4.3.1).
Appendix F Stochastic Length
F.1 Subsequence Selection
In Equation 4, we select a subsequence of length $L$ from the full user history in order to increase sparsity. Our empirical results indicate that careful design of the subsequence selection technique can improve model quality. We compute a metric $f_{i}=t_{n}-t_{i}$ which corresponds to the amount of time elapsed since the user interacted with item $x_{i}$. We conduct offline experiments with the following subsequence selection methods:
  • Greedy Selection – Selects $L$ items with smallest values of $f_{i}$ from $S$
  • Random Selection – Selects $L$ items from $S$ randomly
  • Feature-Weighted Selection – Selects $L$ items from $S$ according to a weighted distribution $1-f_{n,i}/(\sum_{j=1}^{L}f_{j,i})$
During our offline experiments, the feature-weighted subsequence selection method resulted in the best model quality, as shown in Table 13.
Metric Name Selection Type
@@ F.1 Subsequence Selection
Main Engagement Metric (NE) 0.495 0.494 0.495
Main Consumption Metric (NE) 0.792 0.789 0.791
Table 13: Comparison of subsequence selection methods for Stochastic Length on model quality, measured by Normalized Entropy (NE).
F.2 Impact of Stochastic Length on Sequence Sparsity
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
F.3 Comparisons Against Sequence Length Extrapolation Techniques
We conduct additional studies to verify that Stochastic Length is competitive against existing techniques for sequence length extrapolation used in language modeling. Many existing methods perform sequence length extrapolation through modifications of RoPE (Su et al., 2023). To compare against existing methods, we train an HSTU variant (HSTU-RoPE) with no relative attention bias and rotary embeddings.
@@ F.3 Comparisons Against Sequence Length Extrapolation Techniques
Appendix G Sparse Grouped GEMMs and Fused Relative Attention Bias
We provide additional information about the efficient HSTU attention kernel that was introduced in Section 3.2. Our approach builds upon Memory-efficient Attention (Rabe & Staats, 2021) and FlashAttention (Dao et al., 2022), and is a memory-efficient self-attention mechanism that divides the input into blocks and avoids materializing the large $h\times N\times N$ intermediate attention tensors for the backward pass. By exploiting the sparsity of input sequences, we can reformulate the attention computation as a group of back-to-back GEMMs with different shapes. We implement efficient GPU kernels to accelerate this computation. The construction of the relative attention bias is also a bottleneck due to memory accesses. To address this issue, we have fused the relative bias construction and the grouped GEMMs into a single GPU kernel and managed to accumulate gradients using GPU’s fast shared memory in the backward pass. Although our algorithm requires recomputing attention and relative bias in the backward pass, it is significantly faster and uses less memory than the standard approach used in Transformers.
Appendix H Microbatched-Fast Attention Leveraging Cacheable OperatioNs (M-FALCON)
In this section, we provide a detailed description of the M-FALCON algorithm discussed in Section 3.4. We give pseudocode for M-FALCON in Algorithm 1. M-FALCON introduces three key ideas.
(a) GR’s ranking model training (with $n=2n_{c}$ tokens), in causal autoregressive settings.
(b) GR’s ranking model inference utilizing the M-FALCON algorithm.
Figure 11: Illustration of the M-FALCON algorithm. Top: model training in GR’s target-aware formulation. Bottom: model inference with $m$ candidates $\Phi^{\prime}_{0},\ldots,\Phi^{\prime}_{m-1}$, divided into $\lceil m/b_{m}\rceil$ microbatches, where we show model inference for the first microbatch $\Phi^{\prime}_{0},\ldots,\Phi^{\prime}_{b_{m}-1}$ (with $2n_{c}+b_{m}$ total tokens after $\Phi_{0},a_{0},\ldots,\Phi_{n_{c}-1},a_{n_{c}-1}$ are taken into account) above the dotted line. Note that the self-attention algorithm is modified such that $\Phi^{\prime}_{i}$ cannot attend to $\Phi^{\prime}_{j}$ when $i\neq j$ – this is highlighted with “$\times$” in the figure.
Batched inference can be applied to causal autoregressive settings. The ranking task in GR is formulated in a target aware fashion as discussed Section 2.2. Common wisdom suggests that in a target-aware setting, we need to perform inference for one item at a time, with a cost of $O(mn^{2}d)$ for $m$ candidates and a sequence length of $n$. Here we show that this is not the optimal solution; even with vanilla Transformers, we can modify the attention mask used in self-attention to batch such operations (“batched inference”) and reduce cost to $O((n+m)^{2}d)=O(n^{2}d)$.
An illustration is provided in Figure 11. Here, both Figure 11 (a) and (b) involve an attention mask matrix for causal autoregressive settings. The key difference is that Figure 11 (a) uses a standard lower triangular matrix of size $2n_{c}$ for causal training, whereas Figure 11 (b) modifies a lower triangular matrix of size $2n_{c}+b_{m}$ by setting entries for $(i,j)$s where $i,j\geq 2n_{c},i\neq j$ to False or $\infty$- ∞ to prevent target positions $\Phi^{\prime}_{0},\ldots,\Phi^{\prime}_{b_{m}-1}$ from attending to each other. It is easy to see that by doing so, the output of the self-attention block for $\Phi^{\prime}_{i}$, $a^{\prime}_{i}$, only depends on $\Phi_{0},a_{0},\ldots,\Phi_{n_{c}-1},a_{n_{c}-1}$, but not on $\Phi^{\prime}_{j}$ ($i\neq j$). In other words, by making a forward pass over $(2n_{c}+b_{m})$ tokens using the modified attention mask, we can now obtain the same results for the last $b_{m}$ tokens as if we’ve made $b_{m}$ separate forward passes over $(2n_{c}+1)$ tokens, with $\Phi^{\prime}_{i}$ placed at the $2n_{c}$-th (0-based) position during the $i$-th forward pass utilizing a standard causal attention mask.
Microbatching scales batched inference to large candidate sets. Ranking stage may need to deal with a large number of ranking candidates, up to tens of thousands (Wang et al., 2020). We can divide the overall $m$ candidates into $\lceil m/b_{m}\rceil$ microbatches of size $b_{m}$ such that $O(b_{m})=O(n)$, which retains the $O((n+m)^{2}d)=O(n^{2}d)$ running time previously discussed for most practical recommender settings, up to tens of thousands of candidates.
Encoder-level caching enables compute sharing within and across requests. Finally, KV caching (Pope et al., 2022) can be applied both within and across requests. For instance, for the HSTU model presented in this work (Section 3), $K(X)$ and $V(X)$ are fully cachable across microbatches within and/or across requests. For a cached forward pass, we only need to compute $U(X)$, $Q(X)$, $K(X)$, and $V(X)$ for the last $b_{m}$ tokens, while reusing cached $K(X)$ and $V(X)$ for the sequentialized user history containing $n$ tokens. $f_{2}(\text{Norm}{(A(X)V(X))}\odot U(X))$ similarly only needs to be recomputed for the $b_{m}$ candidates. This reduces the cached forward pass’s computational complexity to $O(b_{m}d^{2}+b_{m}nd)$, which significantly improves upon $O((n+b_{m})d^{2}+(n+b_{m})^{2}d)$ by a factor of 2-4 even when $b_{m}=n$.
Algorithm 1 M-FALCON Algorithm.
1:  Input: Merged token series $x_{0},x_{1},\ldots,x_{n-1}$ (can be e.g., $(\Phi_{0},a_{0},\ldots,\Phi_{n_{c}-1},a_{n_{c}-1})$ where $n=2n_{c}$); $m$ ranking candidates $\Phi^{\prime}_{0},\ldots,\Phi^{\prime}_{m-1}$; a $b$-layer $h$-heads self-attention model trained in causal autoregressive settings (e.g., HSTU or Transformers) $f(X,cacheStates,attnMask)\rightarrow(X^{\prime},updatedCacheStates)$ where $X,X^{\prime}\in\mathbb{R}^{N\times d}$, $attnMask\in\mathbb{R}^{N\times N}$, and $cachedStates,updatedCacheStates\in\mathbb{R}^{b\times h\times N\times d_{qk}}\times\mathbb{R}^{b\times h\times N\times d_{qk}}$ (due to caching $K(X)$s and $V(X)$s across $b$ layers); microbatch size $b_{m}$, where we assume $m$ is a multiple of $b_{m}$ for simplicity.
2:  Output: Predictions for all $m$ ranking candidates, $(a^{\prime}_{0},\ldots,a^{\prime}_{m-1})$.
3:  $numMicrobatches=(m+b_{m}-1)//b_{m}$
4:  $attnMask=L_{n+b_{m}}$$L_{n+b_{m}}$ represents a lower triangular matrix. Lower triangular entries are 0s, the rest are $\infty$- ∞.}
5:  $attnMask[i,j]=-\infty$ for $i,j\geq n,i\neq j$This prevents the last $b_{m}$ entries from attending to each other.}
6:  $(a^{\prime}_{0},a^{\prime}_{1},\ldots,a^{\prime}_{b_{m}-1}),kvCache\leftarrow f(embLayer((x_{0},x_{1},\ldots,x_{n-1},\Phi^{\prime}_{0},\ldots,\Phi^{\prime}_{b_{m}-1})),\varnothing,attnMask)$
7:  $predictions=(a^{\prime}_{0},a^{\prime}_{1},\ldots,a^{\prime}_{b_{m}-1})$
8:  $i=1$
9:  while $i<numMicrobatches$ do
10:     $(a^{\prime}_{b_{m}i},a^{\prime}_{b_{m}i+1},a^{\prime}_{b_{m}(i+1)-1}),\_\leftarrow f(embLayer((x_{0},x_{1},\ldots,x_{n-1},\Phi^{\prime}_{b_{m}i},\ldots,\Phi^{\prime}_{b_{m}(i+1)-1})),kvCache,attnMask)$
11:     $predictions\leftarrow predictions+(a^{\prime}_{b_{m}i},a^{\prime}_{b_{m}i+1},\ldots,a^{\prime}_{b_{m}(i+1)-1})$
12:     $i\leftarrow i+1$
13:  end while
14:  return $predictions$
Algorithm 1 is illustrated in Figure 11 to help with understanding. We remark that M-FALCON is not only applicable to HSTUs and GRs, but also broadly applicable as an inference optimization algorithm for other target-aware causal autoregressive settings based on self-attention architectures.
H.1 Evaluation of Inference Throughput: Generative Recommenders (GRs) w/ M-FALCON vs DLRMs
As discussed in Section 3.4, M-FALCON handles $b_{m}$ candidates in parallel to amortize computation costs across all $m$ candidates at inference time. To understand our design, we compare the throughput (i.e., the number of candidates scored per second, QPS) of GRs and DLRMs based on the same hardware setups.
@@ H.1 Evaluation of Inference Throughput: Generative Recommenders (GRs) w/ M-FALCON vs DLRMs
As shown in Figure 12 and Figure 13, GRs’ throughput scales in a sublinear way based on the number of ranking-stage candidates ($m$), up to a certain region – $m=2048$ in our case study – due to batched inference enabling cost amortization. This confirms the criticality of batched inference in causal autoregressive settings. Due to attention complexity scaling as $O((n+b_{m})^{2})$, leveraging multiple microbatches by itself improves throughput. Caching further eliminates redundant linear and attention computations on top of microbatching. The two combined resulted in up to 1.99x additional speedups relative to the $b_{m}=m=1024$ baseline using a single microbatch, as shown in Figure 13. Overall, with the efficient HSTU encoder design and utilizing M-FALCON, HSTU-based Generative Recommenders outperform DLRMs in terms of throughput on a large-scale production setup by up to 2.99x, despite GRs being 285x more complex in terms of FLOPs.
@@
Figure 12: End-to-end inference throughtput: DLRMs vs GRs (w/ M-FALCON) in large-scale industrial settings. Note that this figure is the same as Figure 6, and is reproduced here to facilitate reading.
@@
Figure 13: End-to-end inference throughtput: M-FALCON throughput scaling, on top of the 285x FLOPs GR model, in large batch settings where $m$ (total number of ranking candidates) ranges from 1024 to 16384, and $b_{m}=1024$.